OpenStudy (keana):

Solve each of the quadratic equations below. Describe what the solution represent to the graph of each. x^2 − 25 = 0 and x^2 = 2x + 15

OpenStudy (anonymous):

1) \[x^2-25=0\] \[(x+5)(x-5)=0\] \[x=-5, 5\] This means that the parabola will have 2 x-intercepts at x= -5 and x=5 2) \[x^2=2x+15\] \[x^2-2x-15=0\] \[(x-5)(x+3)=0\] \[x=5,-3\] This means that the parabola will have 2 x-intercepts at x=5 and x=-3 Hope this helps :)